what percent of 240 is 180
Found 2 solutions by stanbon, umail08:
Answer by stanbon(75887)

what percent of 240 is 180
---
Solve
x/100 = 180/240
---
x = 100(3/4)
---
x = 75%
=====================
Cheers,
Stan H.

Answer by umail08(87)

p = percent
b = base
r= rate
p = r/b
p = 180/240
p= 0.75 = 75%